OpenStudy (phebe):

16. Simplify: b. 4√10-√10 @yahman

OpenStudy (phebe):

BRB k XDXDXD

OpenStudy (anonymous):

kk

OpenStudy (phebe):

im back

OpenStudy (anonymous):

4(10−10

OpenStudy (anonymous):

same thing

OpenStudy (astrophysics):

\[4\sqrt{10}-\sqrt{10} \implies 4\sqrt{10}-1\sqrt{10} \implies 3\sqrt{10}\]

OpenStudy (phebe):

thanx guyss